# Learning Discrete Concepts in Latent Hierarchical Models

### v1

@@ 7 Conclusion
: Step 1 shows the availability of the mixture oracle MixOracle (i.e., $|\Omega|$, ${\mathbb{P}}\left(\mathbf{d}=k\right)$, and ${\mathbb{P}}\left(\mathbf{x}|\mathbf{d}=k\right)$ ) as defined in Definition A1. Now, all conditions employed in Theorem A2 are ready, namely Condition 3.1, Condition 3.3 iii, and MixOracle (the consequence of step 1). The derivation in Kivva et al. [24] entails identifying a map from the discrete subspace state index $\mathbf{d}=k$ where $k\in\Omega^{(d)}$ to all discrete components’ state indices $[d_{1},\dots,d_{n_{d}}]=[k_{1},\dots,k_{n_{d}}]$ where $k_{i}\in\Omega^{(d)}_{i}$ is the state index of the $i$-th component $d_{i}$. Thus, we can utilize this map to identify the state index for each individual discrete variable $d_{i}$ from the global index $k$.
@@
To familiarize the reader with the context, we first introduce Theorem A4 and the involved graphical definitions treks A1, t-separation A2, and its connection between d-separation [77].
Definition A1 (Treks).
A trek $T_{i,j}$ in a DAG from vertex $i$ to $j$ consists of a directed path $P_{ki}$ from vertex $k$ to $i$ and a direct path $P_{kj}$ from vertex $k$ to $j$, where we refer to $P_{ki}$ as the $i$ side and $P_{kj}$ as the $j$ side.
@@
To this end, we will introduce a classic theorem (Theorem A6) that connects the non-negative rank of a joint probability table with latent variable states.
Definition A5 (Non-negative Rank).
The non-negative rank of a non-negative matrix $A\in\mathbb{R}^{m\times n}$ is equal to the smallest number $p$ such there exists a non-negative $m\times p$-matrix $\mathbf{B}$ and a non-negative $p\times n$-matrix $\mathbf{C}$ such that $\mathbf{A}=\mathbf{B}\mathbf{C}$.
@@
Given this machinery, we now derive Theorem 3.5 which provides equivalent information in discrete models as Theorem A4 in linear models. See 3.5
@@
We note that $\mathbf{A}\setminus\mathbf{L}$, $\mathbf{B}\setminus\mathbf{L}$, and $\mathbf{L}$ are disjoint because if $\mathbf{A}\cap\mathbf{B}$ is nonempty, it must be a subset of $\mathbf{L}$. Since the graph $\mathcal{G}$ is non-degenerate (Condition 3.1-i) and faithful (Condition 3.7-i), Equation 5 implies the graphical condition that $\mathbf{A}\setminus\mathbf{L}$ and $\mathbf{B}\setminus\mathbf{L}$ are d-separate given $\mathbf{L}$.
@@
Definition A9 (Minimal-graph Operator [19, 20]).
We can merge atomic covers $\mathbf{L}$ into $\mathbf{P}$ in $\mathcal{G}$ if (i) $\mathbf{L}$ is a pure child of $\mathbf{P}$, (ii) all elements of $L$ and $P$ are latent and $|\text{Supp}(\mathbf{L})|=|\text{Supp}(P)|$, and (iii) the pure children of $\mathbf{L}$ form a single atomic cover, or the siblings of $\mathbf{L}$ form a single atomic cover. We denote such an operator as the minimal-graph operator $\mathcal{O}_{\min}(\mathcal{G})$.
This operator merges certain structural redundancies not detectable from rank information [19, 20] (Lemma A13). Please refer to Figure A1 for an example.
@@
Suppose the $\mathcal{G}$ is a linear latent causal model (Definition A7) that satisfies Condition A11. Then the hierarchical causal model $\mathcal{G}$ is identifiable up to the Markov equivalent class of $\mathcal{O}_{s}(\mathcal{O}_{\min}(\mathcal{G}))$.
We note that linear model conditions (Condition A11) and discrete model conditions (Condition 3.7) differ mainly in the substitutes of variables in the linear models with states in the discrete models. This originates from the local graph structures we can access, i.e., states in Theorem 3.5 and variables in Theorem A4. The skeleton operator $\mathcal{O}_{\min}$ (Definition A12 is not necessary under Condition 3.7 since each cover represents a discrete variable whose states must all be connected to its neighbors.
We now present Theorem 3.8 and its proof. See 3.7 See 3.8
@@
We observe that the linearity condition (Definition A7) in Theorem A14 is only utilized to invoke Theorem A4 to access the cardinality of the smallest t-separation set between any two sets of observed variables in the linear model. Through this, the graph identification results in Theorem A14 are derived based on a graph search algorithm repeatedly querying partial graph structures under Condition A11.
For discrete models (Condition 3.1), Theorem 3.5 supplies partial graph structures equivalent to Theorem A4. The difference is that Theorem A4 returns the number of variables in the smallest t-separation set while Theorem 3.5 returns the number of states in the smallest t-separation set. Thus, running Algorithm 1 up to Step 1 (i.e., the original search algorithm Dong et al. [20] with a different rank oracle in Theorem 3.5 highlighted in blue) will return a graph with latent nodes representing discrete states. Algorithm 1 is guaranteed to correctly discover all the atomic covers (Theorem A14) and each atomic cover corresponds to a latent discrete variables (Condition 3.7-ii). Thus, we can obtain each true latent variable by merging all the latent nodes $\mathbf{A}_{L}$ in each atomic cover $\mathbf{A}$ into a discrete latent variable $z$ whose support cardinality $\left\lvert\text{Supp}(z)\right\rvert$ equals to the number of latent nodes $\left\lvert\mathbf{A}_{L}\right\rvert$. We highlight this procedure (Step 1 in Algorithm 1). Moreover, as all latent nodes (i.e., latent states) in an atomic cover belong to one discrete variable, these latent nodes in adjacent atomic covers must be fully connected. Thus, we do not need the skeleton operator $\mathcal{O}_{s}$ as for linear models (Theorem A14). This concludes our proof for Theorem 3.8.
Theorem 3.9 follows the same reasoning as in Theorem 3.8, with the main difference in organizing latent nodes/states into latent discrete variables.
Condition A15 (Discrete Hierarchical Model Conditions for Identifical Supports).
@@
See 3.9
@@
The bulk of the proof overlaps with the proof of Theorem 3.8. Following the same reasoning of the proof of Theorem 3.8, we can obtain a graph with latent nodes representing discrete states before Step 1 and Step 1 in Algorithm 1. Under the identical support condition in Theorem 3.9, we can directly group every $K$ states in an atomic cover into a latent variable as in Algorithm 1-Step 1. Since the true latent variable cardinality is known to be identical, we don’t need Condition 3.1-iiiii to ensure the structure is well defined. Under Condition A15, each atomic cover may contain multiple discrete latent variables, depending on the cover size. It could be possible that one latent variable is not connected to all latent variables in an adjacent atomic cover, as in the linear model case. However, this difference cannot be detected from the rank information (Lemma A13). Thus, we need to retain the skeleton operator $\mathcal{O}_{s}$ inherited from Theorem A14 This concludes our proof for Theorem 3.9. ∎
@@
Figure A1: The discrete graph $\mathcal{G}$ satisfies conditions in Theorem 3.9 (i.e., identical supports). After applying the minimal-graph operator to the graph $\mathcal{G}$, $z_{4}$ is merged to its parent $z_{5}$, and the rank constraints do not change. After applying the skeleton operator to the graph in (b), $z_{1}$ has an edge to $d_{7}$ and $z_{3}$ has an edge to $d_{1}$. We adopt this example from Huang et al. [19].
@@ 9      Orient remaining causal directions that can be inferred from v structures;
11       Theorem 3.9: convert each cover $\mathbf{A}$ to $\log_{K}(\left\lvert\mathbf{A}_{L}\right\rvert)$ discrete variables of cardinality $K$ update $\mathcal{G}^{\prime}$ ;
@@
Algorithm 1 The overall procedure for Rank-based Discrete Latent Causal Model Discovery. [20] We denote the latent nodes in an atomic cover $\mathbf{A}$ as $\mathbf{A}_{L}$ and all observed nodes in the model $\mathcal{G}$ as $\mathbf{X}_{\mathcal{G}}$. We use blue color to highlight our modifications needed for Theorem 3.8 and Theorem 3.9 respectively.
@@ 5             $\mathcal{G^{\prime}}=\text{FindCausalClusters}(\hat{\mathcal{G}},\mathbf{X}_{\mathcal{G}})$;
For the hierarchical model $\mathcal{G}$, we randomly sample the parameters for each causal module, i.e., conditional distributions $p(z_{i}|\text{Parents}(z_{i}))$, according to a Dirichlet distribution over the states of $z_{i}$ with coefficient $1$. For simplicity, we follow condition in Theorem 3.9 and set the support size of latent variables to $2$. Like Kivva et al. [24], we build the generating process from $\mathbf{d}$ to the observed variables $\mathbf{x}$ (i.e., graph $\Gamma$) by a Gaussian mixture model where each state of the discrete subspace corresponds to one component/mode in the mixture model. We truncate the support of each component to improve the invertibility (Condition 3.3-ii). The graphs are exhibited in Figure A2 and Figure A3.
@@
Our method comprise two stages: 1) learning the bottom-level discrete variable $\mathbf{d}$ and the bipartite graph $\Gamma$ from the observed variable $\mathbf{x}$; 2) learning the latent hierarchical model $\mathcal{G}$ given the bottom-level discrete variable $\mathbf{d}$ discovered in 1). For stage 1), we follow the clustering implementation in Kivva et al. [24] under the same hyper-parameter setup as in the original implementation. For stage 2), we apply Algorithm 1 to learn the hierarchical model $\mathcal{G}$. We opt for Step 1 in Algorithm 1 because we evaluate graphs with binary latent variables that meet the conditions of Theorem 3.9. Following Anandkumar et al. [41], Mazaheri et al. [42], we perform conventional rank computation rather than non-negative rank computation and find this replacement satisfactory. We conduct our experiments on a cluster of 64 CPUs. All experiments can be finished within half of an hour. The search algorithm implementation is adapted from Dong et al. [20].
@@
Table 1 and Table 2 correspond to Figure A2 and Figure A3 respectively. As mentioned above, the graphs meet the conditions of Theorem 3.9 with the latent variable cardinality equal to two (binary variables).
@@
For experiments in Section 6.1, we inject concepts by appending keywords to the original prompt. For instance, we inject the concept pair (“sketch”, “wide eyes”) in Figure 4 as follows. For the inference steps $0-10$, we feed the text prompt “A picture of a person”, for steps $10-20$, “a photo of a person, in a sketch style”, and for steps $20-50$, “a photo of a person, in a sketch style, with wide eyes”. For the reverse injection order (injecting “wide eyes” before “sketch”), we inject the following prompts at the three-step stages: “A picture of a person”, “a photo of a person, with wide eyes”, and “a photo of a person, with wide eyes, in a sketch style".
@@
For the attention sparsity experiment (Figure A4), we randomly generate images with the pre-train latent diffusion model and record its attention score across layers. To compute the relative sparsity, we select the threshold as $1/4096$ and compute the proportion of the attention scores over this threshold. For the attention visualization, we randomly select a head from the last attention module in the UNet architecture.
We follow the implementation of Gandikota et al. [55] to train concept sliders of various ranks. We adopt their evaluation protocol to obtain CLIP and LPIPS scores over $20$ randomly sampled images for each rank, concept, and scale combination. We evaluate ranks in $\{2,4,8\}$ and scales $\{1,2,3,4,5\}$. The rank selection technique, inspired by Ding et al. [57], involves multiplying each LoRA’s inner dimension with a scalar parameter and impose $\ell_{0}$ penalty on these scalar parameters. The weight on the $\ell_{0}$ penalty is selected from $\{1e-1,1e-2,1e-3,1e-4,1e-5\}$. We repeat each run for at least three random seeds.
We conduct all our experiments on 2 Nvidia L40 GPUs. Each image inference tasks the same time as in standard SD v1.4 for within two minutes. Each concept slider in Figure 7 takes around half an hour to train.
@@
A5.3 More Examples

### v2

@@ Acknowledgments.
: Step 1 shows the availability of the mixture oracle MixOracle (i.e., $|\Omega|$, ${\mathbb{P}}\left(\mathbf{d}=k\right)$, and ${\mathbb{P}}\left(\mathbf{x}|\mathbf{d}=k\right)$ ) as defined in Definition A1. Now, all conditions employed in Theorem A2 are ready, namely Condition 4.1, Condition 4.3 iii, and MixOracle (the consequence of step 1). The derivation in Kivva et al. [24] entails identifying a map from the discrete subspace state index $\mathbf{d}=k$ where $k\in\Omega^{(d)}$ to all discrete components’ state indices $[d_{1},\dots,d_{n_{d}}]=[k_{1},\dots,k_{n_{d}}]$ where $k_{i}\in\Omega^{(d)}_{i}$ is the state index of the $i$-th component $d_{i}$. Thus, we can utilize this map to identify the state index for each individual discrete variable $d_{i}$ from the global index $k$.
@@
To familiarize the reader with the context, we introduce Theorem A2 and the involved graphical definitions treks 4.6, t-separation 4.7, and its connection between d-separation [84].
See 4.6
@@
To this end, we introduce a classic theorem (Theorem A3) that connects the non-negative rank of a joint probability table with latent variable states. See 4.5
@@
Given this machinery, we now derive Theorem 4.8 which provides equivalent information in discrete models as Theorem A2 in linear models. See 4.8
@@
We note that $\mathbf{A}\setminus\mathbf{L}$, $\mathbf{B}\setminus\mathbf{L}$, and $\mathbf{L}$ are disjoint because if $\mathbf{A}\cap\mathbf{B}$ is nonempty, it must be a subset of $\mathbf{L}$. Since the graph $\mathcal{G}$ is non-degenerate (Condition 4.1-i) and faithful (Condition 4.10-i), Equation 5 implies the graphical condition that $\mathbf{A}\setminus\mathbf{L}$ and $\mathbf{B}\setminus\mathbf{L}$ are d-separate given $\mathbf{L}$.
@@
See 4.11 This operator merges certain structural redundancies not detectable from rank information [19, 20] (Lemma A9). Please refer to Figure A1 for an example.
@@
Suppose the $\mathcal{G}$ is a linear latent causal model (Definition A4) that satisfies Condition A7. Then the hierarchical causal model $\mathcal{G}$ is identifiable up to the Markov equivalent class of $\mathcal{O}_{s}(\mathcal{O}_{\min}(\mathcal{G}))$.
We note that linear model conditions (Condition A7) and discrete model conditions (Condition 4.10) differ mainly in the substitutes of variables in the linear models with states in the discrete models. This originates from the local graph structures we can access, i.e., states in Theorem 4.8 and variables in Theorem A2. The skeleton operator $\mathcal{O}_{\min}$ (Definition A8 is not necessary under Condition 4.10 since each cover represents a discrete variable whose states must all be connected to its neighbors.
We now present Theorem 4.12 and its proof. See 4.10 See 4.12
@@
We observe that the linearity condition (Definition A4) in Theorem A10 is only utilized to invoke Theorem A2 to access the cardinality of the smallest t-separation set between any two sets of observed variables in the linear model. Through this, the graph identification results in Theorem A10 are derived based on a graph search algorithm repeatedly querying partial graph structures under Condition A7.
For discrete models (Condition 4.1), Theorem 4.8 supplies partial graph structures equivalent to Theorem A2. The difference is that Theorem A2 returns the number of variables in the smallest t-separation set while Theorem 4.8 returns the number of states in the smallest t-separation set. Thus, running Algorithm 1 up to Step 1 (i.e., the original search algorithm Dong et al. [20] with a different rank oracle in Theorem 4.8 highlighted in blue) will return a graph with latent nodes representing discrete states. Algorithm 1 is guaranteed to correctly discover all the atomic covers (Theorem A10) and each atomic cover corresponds to a latent discrete variable (Condition 4.10-ii). Thus, we can obtain each true latent variable by merging all the latent nodes $\mathbf{A}_{L}$ in each atomic cover $\mathbf{A}$ into a discrete latent variable $z$ whose support cardinality $\left\lvert\text{Supp}(z)\right\rvert$ equals to the number of latent nodes $\left\lvert\mathbf{A}_{L}\right\rvert$. We highlight this procedure (Step 1 in Algorithm 1). Moreover, as all latent nodes (i.e., latent states) in an atomic cover belong to one discrete variable, these latent nodes in adjacent atomic covers must be fully connected. Thus, we do not need the skeleton operator $\mathcal{O}_{s}$ as for linear models (Theorem A10). This concludes our proof for Theorem 4.12.
Theorem A12 follows the same reasoning as in Theorem 4.12, with the main difference in organizing latent nodes/states into latent discrete variables.
Condition A11 (Discrete Hierarchical Model Conditions for Identical Supports).
@@
We introduce the skeleton operator $\mathcal{O}_{s}$ [19, 20] (Definition A8) that include edges between adjacent covers indistinguishable to rank information.
Theorem A12 (Discrete Hierarchical Identification on Identical Supports).
Suppose the causal model $\mathcal{G}$ satisfies Condition 4.1-i, Condition A11, and $\left\lvert\text{Supp}(z)\right\rvert=K\geq 2$ for all $z\in\mathbf{Z}$. We can identify $\mathcal{G}$ up to the Markov equivalence class of $\mathcal{O}_{s}(\mathcal{O}_{\min}(\mathcal{G}))$.
@@
The bulk of the proof overlaps with the proof of Theorem 4.12. Following the same reasoning of the proof of Theorem 4.12, we can obtain a graph with latent nodes representing discrete states before Step 1 and Step 1 in Algorithm 1. Under the identical support condition in Theorem A12, we can directly group $K$ states in an atomic cover into a latent variable as in Algorithm 1-Step 1. Since the true latent variable cardinality is known to be identical, we don’t need Condition 4.1-iiiii to ensure the structure is well defined. Under Condition A11, each atomic cover may contain multiple discrete latent variables, depending on the cover size. It could be possible that one latent variable is not connected to all latent variables in an adjacent atomic cover, as in the linear model case. However, this difference cannot be detected from the rank information (Lemma A9). Thus, we need to retain the skeleton operator $\mathcal{O}_{s}$ inherited from Theorem A10 This concludes our proof for Theorem A12. ∎
@@
Figure A1: The discrete graph $\mathcal{G}$ satisfies conditions in Theorem A12 (i.e., identical supports). After applying the minimal-graph operator to the graph $\mathcal{G}$, $z_{4}$ is merged to its parent $z_{5}$, and the rank constraints do not change. After applying the skeleton operator to the graph in (b), $z_{1}$ has an edge to $d_{7}$ and $z_{3}$ has an edge to $d_{1}$. We adopt this example from Huang et al. [19].
@@ 9      Orient remaining causal directions that can be inferred from v structures;
11       Theorem A12: convert each cover $\mathbf{A}$ to $\log_{K}(\left\lvert\mathbf{A}_{L}\right\rvert)$ discrete variables of cardinality $K$ update $\mathcal{G}^{\prime}$ ;
@@
Algorithm 1 The overall procedure for Rank-based Discrete Latent Causal Model Discovery. [20] We denote the latent nodes in an atomic cover $\mathbf{A}$ as $\mathbf{A}_{L}$ and all observed nodes in the model $\mathcal{G}$ as $\mathbf{X}_{\mathcal{G}}$. We use blue color to highlight our modifications needed for Theorem 4.12 and Theorem A12 respectively.
@@ 5             $\mathcal{G^{\prime}}=\text{FindCausalClusters}(\hat{\mathcal{G}},\mathbf{X}_{\mathcal{G}})$;
For the hierarchical model $\mathcal{G}$, we randomly sample the parameters for each causal module, i.e., conditional distributions $p(z_{i}|\text{Parents}(z_{i}))$, according to a Dirichlet distribution over the states of $z_{i}$ with coefficient $1$. For simplicity, we follow conditions in Theorem A12 and set the support size of latent variables to $2$. Like Kivva et al. [24], we build the generating process from $\mathbf{d}$ to the observed variables $\mathbf{x}$ (i.e., graph $\Gamma$) by a Gaussian mixture model where each state of the discrete subspace corresponds to one component/mode in the mixture model. We truncate the support of each component to improve the invertibility (Condition 4.3-ii). The graphs are exhibited in Figure A2 and Figure A3.
@@
Our method comprises two stages: 1) learning the bottom-level discrete variable $\mathbf{d}$ and the bipartite graph $\Gamma$ from the observed variable $\mathbf{x}$; 2) learning the latent hierarchical model $\mathcal{G}$ given the bottom-level discrete variable $\mathbf{d}$ discovered in 1). For stage 1), we follow the clustering implementation in Kivva et al. [24] under the same hyper-parameter setup as in the original implementation. For stage 2), we apply Algorithm 1 to learn the hierarchical model $\mathcal{G}$. We opt for Step 1 in Algorithm 1 because we evaluate graphs with binary latent variables that meet the conditions of Theorem A12. Following Anandkumar et al. [52], Mazaheri et al. [53], we perform conventional rank computation rather than non-negative rank computation and find this replacement satisfactory. We conduct our experiments on a cluster of 64 CPUs. All experiments can be finished within half an hour. The search algorithm implementation is adapted from Dong et al. [20].
@@
Table 1 and Table 2 correspond to Figure A2 and Figure A3 respectively. As mentioned above, the graphs meet the conditions of Theorem A12 with the latent variable cardinality equal to two (binary variables).
@@
For experiments in Section 7.1, we inject concepts by appending keywords to the original prompt. For instance, we inject the concept pair (“sketch”, “wide eyes”) in Figure A5 as follows. For the inference steps $0-10$, we feed the text prompt “A picture of a person”, for steps $10-20$, “a photo of a person, in a sketch style”, and for steps $20-50$, “a photo of a person, in a sketch style, with wide eyes”. For the reverse injection order (injecting “wide eyes” before “sketch”), we inject the following prompts at the three-step stages: “A picture of a person”, “a photo of a person, with wide eyes”, and “a photo of a person, with wide eyes, in a sketch style".
@@
For the attention sparsity experiment (Figure A4), we randomly generate images with the pre-train latent diffusion model and record their attention score across layers. To compute the relative sparsity, we select the threshold as $1/4096$ and compute the proportion of the attention scores over this threshold. For the attention visualization, we randomly select a head from the last attention module in the UNet architecture.
We follow the implementation of Gandikota et al. [65] to train concept sliders of various ranks. We adopt their evaluation protocol to obtain CLIP and LPIPS scores over $20$ randomly sampled images for each rank, concept, and scale combination. We evaluate ranks in $\{2,4,8\}$ and scales $\{1,2,3,4,5\}$. The rank selection technique, inspired by Ding et al. [67], involves multiplying each LoRA’s inner dimension with a scalar parameter and imposing $\ell_{0}$ penalty on these scalar parameters. The weight on the $\ell_{0}$ penalty is selected from $\{1e-1,1e-2,1e-3,1e-4,1e-5\}$. We repeat each run for at least three random seeds. The code can be found here.
We conduct all our experiments on 2 Nvidia L40 GPUs. Each image inference takes the same time as in standard SD v1.4 (i.e., within two minutes). Each concept slider in Figure A7 takes around half an hour to train.
@@
A5.3 Discovering Hierarchical Orders from Diffusion Models
We provide further evidence that latent representations at different diffusion steps correspond to different levels of the hierarchical causal model. We select concept pairs, each with higher-level and lower-level concepts. For example, in (“sketch,” “wide eyes”), “sketch” is more global, while “wide eyes” is more local. We alter the text prompt during diffusion generation for concept injection, appending “in a sketch style” to inject “sketch” (see Appendix A5 for prompts). In Figure A5, global concepts are successfully injected at early diffusion steps and local ones at late steps (top row). Reversing this order fails, as shown in the bottom row. For example, injecting “sketch” early and “wide eyes” late renders both correctly, but the global concept “sketch” is absent under the reverse injection order. This supports our theory that concepts are hierarchically organized, with higher-level concepts related to earlier diffusion steps.
Figure A5: Hierarchical Concept Ordering. We inject concepts of distinct abstraction levels into the generating process at different time steps. In the top row, the concept injection follows the hierarchical order, which renders injected concepts faithfully. The bottom row reverses the hierarchical order and cannot incorporate concepts properly. More examples in Figure A8.
A5.4 Causal Sparsity for Concept Extraction
Figure A6 shows that indeed concepts at different abstraction levels have desirable representations at different ranks. For instance, the concept of bright weather is appropriately conveyed by a rank-$2$ LoRA and higher-rank LoRAs alter the background. The same observation occurs to other concepts, where inadequate ranks fail to capture the concept faithfully and unnecessary ranks entangle the target concept with other attributes.
Figure A7 presents the CLIP and LPIPS evaluation for the baseline and our approach, where the CLIP score evaluates the alignment between the image and the target description and the LPIPS score measures the structure change between the edited image and the original image. We can observe that under the sparsity constraint, our approach attains the highest CLIP score and the lowest LPIPS score when compared with the baselines of several ranks, indicating a higher level of alignment and a lower level of undesirable entanglement.
Figure A6: Concepts have varying levels of sparsity. We show that concepts of various abstraction levels correspond to different sparsity levels. For instance, bright weather is appropriately conveyed by a rank-$2$ LoRA and higher-rank LoRAs alter the background. Inadequate ranks fail to capture the concept faithfully and unnecessary ranks entangle the target concept with other attributes.
(a) CLIP scores.
(b) LPIPS scores.
Figure A7: CLIP/LPIPS evaluation. We evaluate our approach and baselines at individual rank constraints. A high CLIP score is favorable as it indicates semantic alignment. A low LPIPS score is more favorable as it indicates minimal excessive changes. We compare our method “sparse” with the optimal fixed rank setting on each concept. For instance, “castle opt: rank4_scale3” indicates that the optimal setting for the concept “castle” is the LoRA of rank $4$ and scale $3$. With a adaptive rank selection, our approach outperforms or keeps up with the optimal fixed setting across different concepts. We repeat each training over three random seeds.
A5.5 More Examples
Figure A8: More examples for Figure A5.